# 3D-LATTE: Latent Space 3D Editing from Textual Instructions

### v1

@@ Abstract
Despite the recent success of multi-view diffusion models for text/image-based 3D asset generation, instruction-based editing of 3D assets lacks surprisingly far behind the quality of generation models. The main reason is that recent approaches using 2D priors suffer from view-inconsistent editing signals. Going beyond 2D prior distillation methods and multi-view editing strategies, we propose a training-free editing method that operates within the latent space of a native 3D diffusion model, allowing us to directly manipulate 3D geometry. We guide the edit synthesis by blending 3D attention maps from the generation with the source object. Coupled with geometry-aware regularization guidance, a spectral modulation strategy in the Fourier domain and a refinement step for 3D enhancement, our method outperforms previous 3D editing methods enabling high-fidelity, precise, and robust edits across a wide range of shapes and semantic manipulations.
Figure 1: 3D-LATTE takes a 3D asset and a user-specified edit instruction as input and leverages the expressiveness of the latent space of a 3D diffusion model to generate edited assets with high-quality details, geometric 3D consistency and precise adherence to diverse edit instructions.
@@ 1 Introduction
Advances in 2D diffusion models Rombach et al. (2021) and 3D representations, such as Neural Radiance Fields (NeRFs) Mildenhall et al. (2020) and 3D Gaussian Splatting (3DGS) Kerbl et al. (2023), have revolutionized 3D asset creation. More recently, powerful text-to-3D Lin et al. (2025); Xiang et al. (2025) and image-to-3D Lin et al. (2025); Xiang et al. (2025); Liu et al. (2023, 2024a) generative models have emerged, enabling scalable and expressive 3D content creation by learning from large 3D data collections Deitke et al. (2023b, a). A core challenge in this domain is instruction-based 3D editing: modifying the geometry and appearance of a 3D object based on a language instruction, while selectively preserving the object’s identity and structure. Achieving semantically accurate, multi-view consistent and high-fidelity edits has attracted considerable attention due to its importance for applications in design, virtual and augmented reality and the entertainment industry, where fine-grained and controllable 3D asset editing is essential.
@@
Many methods propose to solve this task by distilling 2D diffusion priors Brooks et al. (2023) into a 3D representation via score distillation losses Poole et al. (2023); Li et al. (2023); Sella et al. (2023); Koo et al. (2024) or iterative dataset updates Haque et al. (2023); Chen et al. (2023); Chen and Wang (2024). These approaches, however, are constrained by their reliance on 2D supervision. Specifically, they often exhibit multi-view inconsistencies, such as multi-face Janus issues, due to the limited 3D awareness, or inherit 2D editing failures from specific viewpoints, resulting in implausible 3D reconstructions. As a result, most existing methods succeed in appearance edits but struggle with large spatial or geometrical transformations that require globally consistent shape changes. A new line of work Erkoç et al. (2024) proposes to synchronously edit multiple views and relies on feedforward 3D reconstruction models Zhuang et al. (2025); Chen et al. (2024a); Hong et al. (2023) to consolidate them in 3D or adopts a hybrid 2D-3D approach Chen et al. (2024b), where multi-view images are fused into a 3D representation at each denoising step. Nevertheless, existing methods based on feedforward 3D reconstruction often suffer from blurry, distorted reconstructions, due to inconsistent edits across views being propagated to the 3D model. Hybrid 2D-3D methods, on the other hand, may introduce Janus artifacts and multi-view inconsistencies due to the use of 2D priors at intermediate steps.
@@
To overcome the aforementioned limitations, we take an alternative approach and leverage a 3D-native diffusion prior, where noise is injected directly to the 3D representation. This grants us the ability to directly manipulate the appearance and geometry of a 3D asset without reliance on 2D priors and costly optimization procedures. Importantly, we operate within the model’s latent space, which is structured as pixel-aligned 3D Gaussian representations Lin et al. (2025). Motivated by the role of attention maps in 2D editing Hertz et al. (2024), our key insight is that the 3D self- and cross-attention maps naturally capture information about the layout and composition of the 3D scene and directly model relationships between 3D Gaussians and text tokens. Building on this observation, we introduce a 3D attention injection mechanism into the denoising process of a text-conditioned 3D diffusion model. At each time step, we modulate the model’s attention maps by blending or replacing them with those obtained from the generation with the source prompt, namely the description of the original, unedited asset. This allows us to guide the model to synthesize a 3D asset that semantically aligns with the edit prompt while preserving the structural composition and geometry of the source asset.
@@
To achieve localized and controlled edits, we leverage a vision-language model (VLM) in combination with a segmentation model to generate multi-view consistent 2D masks. These masks naturally define a 3D segmentation over the multi-view pixel-aligned 3D Gaussians, allowing us to constrain the modification to the relevant 3D regions. To enhance 3D quality and robustness, we adopt a frequency-modulated strategy that emphasizes low-frequency components early in the denoising process, encouraging the model to capture global structure before refining fine-grained details. Structural coherence is further reinforced through a geometry-aware regularization term, applied in the form of classifier guidance. Finally, to address edits requiring higher fidelity outputs we adopt an iterative strategy that progressively enhances the fine-grained details of the 3D representation while preserving cross-view appearance consistency.
@@
To demonstrate the effectiveness of our approach, we conduct extensive user studies, report quantitative metrics based on CLIP similarity Haque et al. (2023) and evaluate performance using GPTEval3D Wu et al. (2024b), consistently surpassing previous state-of-the-art works.
@@ 2 Related Works
Recent methods in 3D generation have shifted their focus toward diffusion-based models Ho et al. (2020), which have been successfully adapted to a plethora of 3D representations, including voxel grids Hui et al. (2022), point clouds Zeng et al. (2022); Nichol et al. (2022) and triplanes Shue et al. (2023); Anciukevicius et al. (2022). More recently, methods leveraging 3DGS Yuanbo et al. (2025); Zhou et al. (2024); Lin et al. (2025); Zhang et al. (2024) have emerged as particularly promising. GaussianCube Zhang et al. (2024) organizes a 3DGS representation into a structured voxel field and applies 3D diffusion using a standard 3D U-Net. In contrast, DiffSplat Lin et al. (2025) introduces a 3D latent diffusion model that directly generates 3D Gaussians by modeling an object as a set of multi-view 3DGS grids.
2.2 3D Editing with 2D Priors
@@
Editing in 3D presents unique challenges due to the need for multi-view consistency. One line of work tackles this by leveraging image-conditioned 2D diffusion models such as InstructPix2Pix Brooks et al. (2023), injecting their editing capabilities into learned 3D representations. A pioneering method in this direction, InstructNeRF2NeRF Haque et al. (2023), introduces the “Iterative Dataset Update” method, where rendered views from the 3D model are repeatedly updated during optimization. More recent methods, such as GaussianEditor Chen et al. (2023) and ProEdit Chen and Wang (2024), build upon this paradigm. However, large edits can still lead to multi-view inconsistencies due to the lack of explicit multi-view awareness during editing. Another line of work, such as DreamEditor Zhuang et al. (2023), FocalDreamer Li et al. (2023) and Vox-E Sella et al. (2023), relies on Score Distillation Sampling (SDS) to guide 3D asset synthesis by distilling gradients from a pre-trained diffusion model. Posterior Distillation Sampling Koo et al. (2024) improves upon SDS by aligning the latents of source and target images. While effective, these methods inherit known limitations of SDS, including oversmoothed textures, Janus artifacts, and mode-seeking behavior. As an alternative, recent methods Wu et al. (2024a); Lee et al. (2025); Chen et al. (2024c) perform multi-view consistent edits directly on source images, which are then used to update the 3D representation. For instance, DGE Chen et al. (2024c) leverages epipolar constraints to aggregate features across views, while GaussCTRL Wu et al. (2024a) performs depth-conditioned 2D updates combined with cross-view alignment. However, such reliance on depth guidance can limit the model’s ability to perform significant shape changes. In contrast, by operating directly within the latent space of a 3D diffusion generative model, our method enables flexible manipulation of both appearance and geometry in a fully 3D-consistent manner without relying on 2D priors or inheriting the limitations of SDS-based optimization.
2.3 Hybrid 2D-3D Editing Methods
@@
In parallel, another line of work has emerged that moves beyond pure 2D updates by incorporating hybrid 2D-3D strategies. SHAP-Editor Chen et al. (2024d) learns a feed-forward editor operating in the latent space of Shap-E Jun and Nichol (2023) and is trained using an SDS Poole et al. (2023) objective. However, it requires retraining for each new set of edit instructions and relies heavily on the 2D priors and latent structure of Shap-E Jun and Nichol (2023). Thus, its edits lack flexibility and are often of poor visual quality, limiting its applicability in in-the-wild scenarios. In contrast, our method generalizes across categories, produces high-fidelity outputs, and enables fast inference. MVEdit Chen et al. (2024b) adopts a hybrid 2D-3D approach by fusing multi-view images into a 3D representation between denoising steps in a multi-view diffusion model. While this is a promising step toward enabling 3D-aware 2D edits, its reliance on 2D priors can still introduce Janus artifacts and multi-view inconsistencies.
3 Methodology
Given a 3D object, a source text prompt $p$ describing its original appearance (e.g., a dog) and a target text prompt $p^{\ast}$ describing the desired edit (e.g., a dog wearing sunglasses), our goal is to alter the object’s appearance and/or geometry so that it semantically aligns with the target prompt. At the same time, we aim to preserve regions not referenced by the prompt and maintain the original structure of the object.
@@ 3 Methodology
To this end, we introduce a zero-shot editing framework that extends attention control to the domain of 3DGS. Our method operates within the text-guided 3D diffusion model DiffSplat Lin et al. (2025), outlined in Section 3.1. Our core idea is to inject cross- and self-attention maps derived from the source 3D asset during the diffusion process that synthesizes the edited 3D asset as explained in Section 3.3. This enables semantically meaningful edits while maintaining multi-view consistency and 3D structural coherence, since our attention modifications are applied within a latent space that encodes 3D geometry. To further enhance quality, we incorporate a geometry-aware regularization mechanism (Section 3.5) and a frequency annealing strategy (Section 3.6). Finally, in Section 3.7 we introduce an iterative refinement pipeline that progressively enhances high-frequency detail and texture fidelity in the reconstructed 3D asset. An overview of our method is illustrated in Figure 2.
@@
Figure 2: Overview of 3D-LATTE. We operate in the latent space of a pre-trained 3D generative model. The source 3D object is represented as a multi-view Gaussian splat grid and inverted into its corresponding noise latent. Starting from this latent, we perform denoising guided by the edit prompt, while injecting 3D cross- and self-attention maps derived from the source object. A geometry regularization guidance term, a frequency modulation strategy and a 3D enhancement module further refine the result. Region-specific edits are supported via masks generated using GroundingDINO Liu et al. (2024b) and SAM2 Ravi et al. (2024).
@@
In this work we leverage DiffSplat Lin et al. (2025) as our 3D diffusion backbone. In Lin et al. (2025) a 3D object is modeled as a set of structured multi-view splat grids $\mathcal{G}=\{G_{i}\}_{i=1}^{V_{\text{in}}}$, where each $G_{i}\in\mathbb{R}^{C\times H\times W}$, $V_{\text{in}}$ is the number of input views, $H\times W$ is the spatial resolution, and $C=12$ corresponds to the number of Gaussian attributes. Each Gaussian primitive $g_{i}\in\mathbb{R}^{12}$ is parameterized by its RGB color $\mathbf{c_{i}}\in\mathbb{R}^{3}$, 3D location $\mathbf{x_{i}}\in\mathbb{R}^{3}$, which is determined by depth $\mathbf{d_{i}}\in\mathbb{R}$ and the camera parameters, scale $\mathbf{s_{i}}\in\mathbb{R}^{3}$, rotation quaternion $\mathbf{r_{i}}\in\mathbb{R}^{4}$, and opacity $\mathbf{o_{i}}\in\mathbb{R}$. The pipeline of Lin et al. (2025) comprises a Gaussian reconstruction module that converts multi-view RGB images, along with depth and normal maps into a Gaussian splat grid representation, a VAE that encodes the Gaussian splat grid into a latent space and a generative model that is trained on this representation. Finally, the reconstructed latent is fed to a VAE decoder to obtain the 3D gaussian representation, which can be rendered as multi-view images.
@@
Following Lin et al. (2025), we represent a 3D object as a set of structured multi-view splat grids $\mathcal{G}=\{G_{i}\}_{i=1}^{V_{\text{in}}}$. We use the reconstruction module from DiffSplat Lin et al. (2025), which takes as input a set of $V_{\text{in}}$ RGB images, depth maps, and normal maps rendered from uniformly distributed camera viewpoints around the source 3D object, resulting in a 3D Gaussian for each pixel of the input views. We begin by inverting the source 3D object representation into its corresponding noise latent $z_{T}$. However, standard DDIM inversion Song et al. (2021) can introduce slight errors, leading to poor reconstructions. Thus, we adapt the DDPM inversion mechanism proposed in Huberman-Spiegelglas et al. (2024) to operate on a set of multi-view Gaussian splat grids $\mathcal{G}$ and recover a noise vector that accurately reconstructs the original 3D object.
Let the encoded Gaussian splat grids (i.e., splat latents) be denoted as $z_{0}=\{z_{0}^{i}\}_{i=1}^{V_{\text{in}}}=\{E_{\phi}(G_{i})\}_{i=1}^{V_{\text{in}}}$ where ${E_{\phi}}(G)$ is the Gaussian VAE encoder of Lin et al. (2025). We construct the forward diffusion trajectory Song et al. (2021):
$z_{t}=\sqrt{\alpha_{t}}\cdot z_{0}+\sqrt{1-\alpha_{t}}\cdot\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,I)$ (1)
where $\alpha_{t}$ is the cumulative diffusion noise schedule and $\epsilon_{t}\sim N(0,I)$ is sampled independently per timestep. To obtain the noise maps, we extract:
$\phi_{t}=(z_{t-1}-\mu_{\theta}(z_{t},t))/{\sigma_{t}}$ (2)
where $\mu_{\theta}$ and $\sigma_{t}$ are the predicted mean and standard deviation of the noise injected at each timestep, respectively. This sequence of noise maps is shown in Huberman-Spiegelglas et al. (2024) to be more edit-friendly since they are constructed with statistically independent noise samples. Starting from $z_{T}$, we perform the denoising process with the edit prompt $p^{\ast}$. To achieve our goal of editing while maintaining 3D structural information and layout, we apply 3D self- and cross-attention injection.
@@
In our formulation, 3D-aware attention refers to the self- and cross-attention maps over the noisy multi-view Gaussian latents. In the diffusion backbone Lin et al. (2025) at each timestep, features of the noisy Gaussian splat latents $\phi({z}_{t})\in\mathbb{R}^{V\times D\times H\times W}$ are projected into a query matrix, $Q$, while the keys, $K$, are obtained from the text prompt embeddings. The cross-attention map (omitting head and batch dimensions for clarity) is computed as:
$A^{\text{cross}}=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)\in\mathbb{R}^{V\times(HW)\times T}$ (3)
where $T$ is the number of text tokens. Each entry $A_{v,i,j}$ represents the influence of the $j^{\text{th}}$ token on the $i^{\text{th}}$ Gaussian latent in view $v$. This cross-attention mechanism injects semantic information into the 3D latent representation during the denoising process, thus guiding splat-to-text interactions.
Self-attention maps capture spatial and semantic relationships between all 3D Gaussians. Both queries and keys are computed via linear projections of the Gaussian features:
$A^{\text{self}}=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)\in\mathbb{R}^{(VHW)\times(VHW)}$ (4)
@@
Motivated by the role of attention in encoding 3D structure and semantics within the 3D Gaussian splat representation, we introduce a 3D attention injection framework. Let $\text{DM}(z_{t},t,p)$ denote the denoising step with the source prompt $p$, which computes $z_{t-1}$, and $\text{DM}(z_{t}^{*},t,p^{*})$ be the denoising step with the edit prompt $p^{*}$, which computes $z^{\ast}_{t-1}$. Additionally, let $A_{t}=(A_{t}^{\text{cross}},A_{t}^{\text{self}})$ and $A_{t}^{\ast}=((A_{t}^{\ast})^{\text{cross}},(A_{t}^{\ast})^{\text{self}})$ be the attention maps computed in the denoising steps $\text{DM}(z_{t},t,p)$ and $\text{DM}(z_{t}^{*},t,p^{*})$, respectively. We inject modified attention maps $\hat{A}_{t}=(\hat{A}_{t}^{\text{cross}},\hat{A}_{t}^{\text{self}})$, derived from the source prompt $p$, into the denoising process guided by the edit prompt $p^{\ast}$, as follows. At each denoising step with prompt $p^{\ast}$, we override the attention maps $A^{\ast}_{t}$ computed in $z_{t-1}^{\ast}=\text{DM}(z_{t}^{\ast},t,p^{\ast})$ with $\hat{A}_{t}$ (i.e., $A^{\ast}_{t}\leftarrow\hat{A}_{t}$). To calculate the modified cross-attention map, we follow Hertz et al. (2024) and inject attention until a timestep $\tau_{\text{cross}}$ only on tokens shared between the original and edited prompts, using the following $Edit$ function:
$\hat{A}_{t}^{\text{cross}}=Edit(A^{\text{cross}},(A^{\ast})^{\text{cross}},t)_{i,j}:=\begin{cases}((A_{t}^{\ast})^{\text{cross}})_{i,j}&\text{if }CT(j)=None\text{ and }t>\tau_{\text{cross}}\\ (A_{t}^{\text{cross}})_{i,CT(j)}&\text{otherwise}\end{cases}$

### v3

@@ Abstract
Despite the recent success of multi-view diffusion models for text/image-based 3D asset generation, instruction-based editing of 3D assets lacks surprisingly far behind the quality of generation models. The main reason is that recent approaches using 2D priors suffer from view-inconsistent editing signals. Going beyond 2D prior distillation methods and multi-view editing strategies, we propose a training-free editing method that operates within the latent space of a native 3D diffusion model, allowing us to directly manipulate 3D geometry. We guide the edit synthesis by blending 3D attention maps from the generation with the source object. Coupled with geometry-aware regularization guidance, a spectral modulation strategy in the Fourier domain and a refinement step for 3D enhancement, our method outperforms previous 3D editing methods enabling high-fidelity and precise edits across a wide range of shapes and semantic manipulations. Our project webpage is: https://mparelli.github.io/3d-latte.
Figure 1: 3D-LATTE takes a 3D asset and a user-specified edit instruction as input and leverages the expressiveness of the latent space of a 3D diffusion model to generate edited 3D assets with high-quality details, geometric 3D consistency and precise adherence to diverse edit instructions.
@@ 1 Introduction
Advances in 2D diffusion models [rombach2021highresolution] and 3D representations, such as Neural Radiance Fields (NeRFs) [mildenhall2020nerf] and 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians], have revolutionized 3D asset creation. More recently, powerful text-to-3D [lin2025diffsplat, xiang2024structured] and image-to-3D [lin2025diffsplat, xiang2024structured, liu2023zero1to3, liu2023one2345++] generative models have emerged, enabling scalable and expressive 3D content creation by learning from large 3D data collections [Deitke2023objaverse, objaverseXL]. A core challenge in this domain is instruction-based 3D editing: modifying the geometry and appearance of a 3D object based on a language instruction, while selectively preserving the object’s identity and structure. Achieving semantically accurate and high-fidelity 3D edits has attracted considerable attention due to its importance for applications in design and virtual and augmented reality.
@@
Many methods propose to solve this task by distilling 2D diffusion priors [brooks2023instructpix2pix] into a 3D representation via score distillation losses [poole2022dreamfusion, li2023focaldreamer, sella2023vox, Koo2024PDS] or iterative dataset updates [Haque2023instructnerf, chen2023gaussianeditor, Chen2024proedit]. These approaches, however, are constrained by their reliance on 2D supervision. Specifically, they often exhibit multi-view inconsistencies, such as multi-face Janus issues, due to the limited 3D awareness, or inherit 2D editing failures from specific viewpoints, resulting in implausible 3D reconstructions. As a result, most existing methods succeed in appearance edits but struggle with large spatial or geometrical transformations that require globally consistent shape changes. A new line of work [erkoc2024preditor3d, li2025cmd, huang2025edit360] proposes to synchronously edit multiple views via multi-view diffusion priors or relies on feedforward 3D reconstruction models [zhuang2024gtr, LaRa, hong2023lrm] to consolidate them in 3D. Others adopt a hybrid 2D-3D approach [mvedit2024], where multi-view images are fused into a 3D representation at each denoising step. Nevertheless, existing methods based on feedforward 3D reconstruction or multi-view priors often suffer from blurry, distorted reconstructions, due to small inconsistencies across views being propagated to the 3D model. Hybrid 2D-3D methods, on the other hand, may introduce Janus artifacts and multi-view inconsistencies due to the use of 2D priors at intermediate steps.
@@
To overcome these limitations, we take an alternative approach and leverage a 3D-native diffusion prior, where noise is injected directly to the 3D representation. This allows us to directly manipulate the appearance and geometry of a 3D asset without reliance on 2D or multi-view priors. Importantly, we operate within the model’s latent space, which is structured as pixel-aligned 3D Gaussian representations [lin2025diffsplat]. Motivated by the role of attention maps in 2D editing [hertz2022prompt], our key insight is that the 3D self- and cross-attention maps naturally capture information about the layout and composition of the 3D scene and directly model relationships between 3D Gaussians and text tokens. Building on this observation, we introduce a 3D attention injection mechanism into the denoising process of a text-conditioned 3D diffusion model. At each time step, we modulate the model’s attention maps by blending or replacing them with those obtained from the generation with the source prompt, namely the description of the original, unedited asset. This allows us to guide the model to synthesize a 3D asset that semantically aligns with the edit prompt while preserving the 3D structure of the source asset.
@@
To achieve localized edits, we leverage a vision-language model (VLM) in combination with a segmentation model to generate multi-view consistent 2D masks. These masks naturally define a 3D segmentation over the multi-view pixel-aligned 3D Gaussians, allowing us to constrain the modification to the relevant 3D regions. To enhance 3D quality, we adopt a frequency-modulated strategy that emphasizes low-frequency components early in the denoising process, encouraging the model to capture global structure before refining fine-grained details. Structural coherence is further reinforced through a geometry-aware regularization term, applied in the form of classifier guidance. Finally, to address higher-fidelity edits we adopt an iterative strategy that progressively enhances the fine-grained details of the 3D representation while preserving cross-view consistency.
@@
To demonstrate the effectiveness of our approach, we conduct extensive user studies, report quantitative metrics based on CLIP similarity [Haque2023instructnerf] and evaluate performance using GPTEval3D [wu2023gpteval3d], consistently surpassing previous state-of-the-art works.
@@ 2 Related Works
3D Diffusion Generative Models. Recent methods in 3D generation have shifted their focus toward diffusion-based models [ho2020denoising], which have been adapted to a plethora of 3D representations, including voxel grids [hui2022neural, trellis], point clouds [zeng2022lion, nichol2022pointe] and triplanes [Shue2023TriDiff, anciukevicius2022renderdiffusion]. Recently, methods leveraging 3DGS [yang2024prometheus, Zhou2025diffgs, lin2025diffsplat, zhang2024gaussiancube] have emerged as particularly promising. GaussianCube [zhang2024gaussiancube] organizes a 3DGS representation into a voxel field and applies 3D diffusion using a 3D U-Net. In contrast, DiffSplat [lin2025diffsplat] introduces a 3D latent diffusion model that directly generates 3D Gaussians by modeling an object as a set of multi-view 3DGS grids.
@@
3D Editing with 2D Priors. Editing in 3D presents unique challenges due to the need for multi-view consistency. One line of work tackles this by leveraging image-conditioned 2D diffusion models such as InstructPix2Pix [brooks2023instructpix2pix], injecting their editing capabilities into learned 3D representations. A pioneering work in this direction, InstructNeRF2NeRF [Haque2023instructnerf], introduces the “Iterative Dataset Update” method, where rendered views from the 3D model are repeatedly updated during optimization. More recent methods [chen2023gaussianeditor, Chen2024proedit], build upon this paradigm. However, large edits can still lead to multi-view inconsistencies due to the lack of explicit multi-view awareness. Another line of work,  [zhuang2023dreameditor, li2023focaldreamer, sella2023vox], relies on Score Distillation Sampling (SDS) to guide 3D asset synthesis by distilling gradients from a pre-trained diffusion model. Posterior Distillation Sampling [Koo2024PDS] improves upon SDS by aligning the latents of source and target images. While effective, these methods inherit known limitations of SDS, including Janus artifacts and mode-seeking behavior. As an alternative, recent methods [wu2024gaussctrl, lee2025editsplat, chen2024dge, wen2025intergsedit] perform multi-view consistent edits directly on source images, which are then used to update the 3D representation. For instance, DGE [chen2024dge] leverages epipolar constraints to aggregate features across views, while GaussCTRL [wu2024gaussctrl] performs depth-conditioned 2D updates combined with cross-view alignment. However, such reliance on depth guidance can prevent the model from performing significant shape changes. In contrast, by operating directly within the latent space of a 3D diffusion model, our method enables flexible manipulation of both appearance and geometry in a fully 3D-consistent manner without relying on 2D editing priors or SDS-based optimization.
@@
Hybrid 2D-3D Editing Methods. In parallel, another line of work has emerged that moves beyond pure 2D updates by incorporating hybrid 2D-3D strategies. SHAP-Editor [chen2024shap] learns a feed-forward editor operating in the latent space of Shap-E [Jun2023ShapEGC] and is trained using an SDS [poole2022dreamfusion] objective. However, it requires retraining for each new set of edits and relies heavily on the 2D priors and latent structure of Shap-E [Jun2023ShapEGC]. Thus, its edits lack flexibility and are often of poor visual quality. In contrast, our method generalizes across categories, produces high-fidelity outputs, and enables fast inference. MVEdit [mvedit2024] adopts a hybrid 2D-3D approach by fusing images into a 3D representation between denoising steps in a multi-view diffusion model. More recently, Edit360 [huang2025edit360] builds on a dense-view synthesis model and achieves propagation of edits by aligning an edited view camera trajectory with the original front-view trajectory. While these are promising steps toward enabling 3D-aware edits, reliance on 2D priors or trajectory alignment strategies can introduce multi-view inconsistencies.
3 Method
@@
Given a 3D object, a source text prompt $p$ describing its original appearance and a target text prompt $p^{\ast}$ describing the desired edit, our goal is to alter the object’s appearance and/or geometry so that it semantically aligns with the target prompt. Simultaneously, we aim to preserve regions not referenced by the prompt and maintain the original object structure. To this end, we introduce a zero-shot editing framework that extends attention control to the domain of 3DGS. Our method operates within the text-guided 3D diffusion model DiffSplat [lin2025diffsplat], outlined in Section 3.1. Our core idea is to inject 3D cross- and self-attention maps derived from the source 3D asset during the diffusion process that synthesizes the edited 3D asset as explained in Section 3.3. This enables semantically meaningful edits while maintaining multi-view consistency and 3D structural coherence, since our attention modifications are applied within a latent space that encodes 3D geometry. To further enhance quality, we incorporate a geometry-aware regularization mechanism (Section 3.5) and a frequency annealing strategy (Section 3.6). Finally, in Section 3.7 we introduce an iterative refinement pipeline that progressively enhances high-frequency details and texture fidelity in the reconstructed 3D asset. An overview of our method is shown in Figure 2.
@@
Figure 2: Overview of 3D-LATTE. We operate in the latent space of a pre-trained 3D generative model. The source 3D object is represented as a multi-view Gaussian splat grid and inverted into its corresponding noise latent. Starting from this latent, we perform denoising guided by the edit prompt, while injecting 3D cross- and self-attention maps derived from the source object. A geometry regularization guidance term, a frequency modulation strategy and a 3D enhancement module further refine the result. Region-specific edits are supported via masks generated using GroundingDINO [liu2023grounding] and SAM2 [ravi2024sam2].
@@
In this work, we leverage DiffSplat [lin2025diffsplat] as our 3D diffusion backbone. In [lin2025diffsplat] a 3D object is modeled as a set of structured multi-view splat grids $\mathcal{G}=\{G_{i}\}_{i=1}^{V}$, where each $G_{i}\in\mathbb{R}^{C\times H\times W}$, $V$ is the number of input views, $H\times W$ is the spatial resolution, and $C=12$ corresponds to the number of Gaussian attributes. Each Gaussian primitive $g_{i}\in\mathbb{R}^{12}$ is parameterized by its RGB color $c_{i}\in\mathbb{R}^{3}$, 3D location $x_{i}\in\mathbb{R}^{3}$, which is determined by its depth and camera parameters, scale $s_{i}\in\mathbb{R}^{3}$, rotation quaternion $r_{i}\in\mathbb{R}^{4}$, and opacity $o_{i}\in\mathbb{R}$. The pipeline of [lin2025diffsplat] comprises a Gaussian reconstruction module that converts multi-view RGB images, along with depth and normal maps into a Gaussian splat grid representation, a VAE that encodes the Gaussian splat grid into a latent space and a generative model that is trained on this representation. Finally, the reconstructed latent is fed to a VAE decoder to obtain the 3D gaussian representation, which can be rendered as multi-view images.
@@
Following [lin2025diffsplat], we represent a 3D object as a set of structured multi-view splat grids $\mathcal{G}=\{G_{i}\}_{i=1}^{V}$. We use the reconstruction module from DiffSplat [lin2025diffsplat], which takes as input a set of $V$ RGB images, depth maps, and normal maps rendered from uniformly distributed camera viewpoints around the source 3D object, resulting in a 3D Gaussian for each pixel of the input views. We begin by inverting the source 3D object representation into its corresponding noise latent $z_{T}$. However, standard DDIM inversion [song2021ddim] can introduce slight errors, leading to poor reconstructions. Thus, we adapt the DDPM inversion mechanism proposed in  [huberman2024edit] to operate on a set of multi-view Gaussian splat grids $\mathcal{G}$ and recover a noise vector that accurately reconstructs the original 3D object.
Let the encoded Gaussian splat grids (i.e., splat latents) be denoted as $z_{0}=\{z_{0}^{i}\}_{i=1}^{V}=\{E_{\phi}(G_{i})\}_{i=1}^{V}$ where ${E_{\phi}}(\cdot)$ is the Gaussian VAE encoder of [lin2025diffsplat]. We construct the forward diffusion trajectory [song2021ddim] and obtain the noise maps $\eta_{t}$ as follows:
$z_{t}=\sqrt{\alpha_{t}}\,z_{0}+\sqrt{1-\alpha_{t}}\,\epsilon_{t},\quad\eta_{t}=\frac{(z_{t-1}-\mu_{\theta}(z_{t},t))}{\sigma_{t}}$ (1)
where $\alpha_{t}$ is the cumulative diffusion noise schedule, $\epsilon_{t}\sim\mathcal{N}(0,I)$ is sampled independently per timestep, and $\mu_{\theta}$ and $\sigma_{t}$ are the predicted mean and standard deviation of the noise injected at each timestep, respectively. This sequence of noise maps is shown in [huberman2024edit] to be edit-friendly since they are constructed with statistically independent noise samples. Starting from $z_{T}$, we perform the denoising process with the edit prompt $p^{\ast}$. To achieve editing while maintaining 3D structural information and layout, we apply 3D self- and cross-attention injection.
@@
In our formulation, 3D-aware attention refers to the self- and cross-attention maps over the noisy multi-view Gaussian latents. In our 3D diffusion backbone at each timestep, features of the noisy Gaussian splat latents $\phi({z}_{t})\in\mathbb{R}^{V\times D\times H\times W}$ are projected into a query matrix, $Q$, while the keys, $K$, and values, $V$, are obtained from the text prompt embeddings. Each entry $W_{i,j}$ in the 3D cross-attention map $W_{\mathcal{G}}^{\text{cross}}$ represents the influence of the $j$-th text token on the $i$-th Gaussian latent. As illustrated in Figure 3, cross-attention weights define a token-3D Gaussian splat correspondence field, enabling accurate 3D localization of interest regions. Self-attention maps $W_{\mathcal{G}}^{\text{self}}$ capture spatial and semantic relationships between all 3D Gaussian latents.
Figure 3: Per-token attention over 3D splats, rendered from multiple viewpoints.
Queries, keys and values are computed via linear projections of the Gaussian features.
@@
Motivated by the role of attention in encoding 3D structure and semantics within the 3D Gaussian splat representation, we introduce a 3D attention injection framework. Let $z_{t-1}=D_{\theta}(z_{t},t,p)$ denote the denoising step with the source prompt $p$, and $z^{\ast}_{t-1}=D_{\theta}(z_{t}^{*},t,p^{*})$ be the denoising step with the edit prompt $p^{*}$. Additionally, let ${W_{\mathcal{G}}}_{t}=({W_{\mathcal{G}}}^{\text{cross}}_{t},{W_{\mathcal{G}}}_{t}^{\text{self}})$ and ${W_{\mathcal{G}}}_{t}^{\ast}=(({W_{\mathcal{G}}}_{t}^{\ast})^{\text{cross}},({W_{\mathcal{G}}}_{t}^{\ast})^{\text{self}})$ be the attention maps computed in the denoising steps $D_{\theta}(z_{t},t,p)$ and $D_{\theta}(z_{t}^{*},t,p^{*})$, respectively. We inject modified attention maps $\hat{W_{\mathcal{G}}}_{t}=(\hat{W_{\mathcal{G}}}_{t}^{\text{cross}},\hat{W_{\mathcal{G}}}_{t}^{\text{self}})$, derived from the source prompt $p$, into the denoising process guided by the edit prompt $p^{\ast}$, as follows. At each denoising step with prompt $p^{\ast}$, we override the attention maps ${W_{\mathcal{G}}}^{\ast}_{t}$ with $\hat{W_{\mathcal{G}}}_{t}$ (i.e., ${W_{\mathcal{G}}}^{\ast}_{t}\leftarrow\hat{W_{\mathcal{G}}}_{t}$). To calculate the modified cross-attention map, we follow [hertz2022prompt] and inject 3D attention until a timestep $\tau_{\text{cross}}$ only on tokens shared between the original and edited prompts, using the following $F_{cross}^{3D}$ function:
$\hat{W_{\mathcal{G}}}_{t}^{\text{cross}}$ $=F_{\text{cross}}^{3D}\!\left(W_{\mathcal{G}}^{\text{cross}},(W_{\mathcal{G}}^{\ast})^{\text{cross}},t\right)_{i,j}$ (2)
[…]
where $CT$ is an alignment function that maps a token index from $p^{\ast}$ to the corresponding token index in $p$. In the case where there is no match, it returns an empty set.
Figure 4: We form the normalized Laplacian of the 3D self-attention graph and map the first three eigenvectors to a color value per Gaussian.